# Time-Series Foundation Model for Value-at-Risk Forecasting

### v5

@@ 8 CRediT author statement
Anubha Goel: Data curation; Formal analysis; Investigation; Validation; Visualization; Writing- Original draft preparation. Puneet Pasricha: Formal analysis; Investigation; Methodology; Validation; Visualization; Writing- Original draft preparation. Juho Kanniainen: Conceptualization; Methodology; Project administration; Supervision; Validation; Visualization; Writing- Original draft preparation; Writing- Reviewing and Editing.
9 Acknowledgements
The first author acknowledges support from the European Union’s Horizon Europe programme under the Marie Skłodowska-Curie Actions (Grant Agreement No. 101150609, Project: HiddenTipChains).
Figure 1: This figure shows the bar plots of number of assets (out of 92) for which we failed to reject the null hypothesis in the Unconditional Coverage (UC) test at various value-at-risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.
Figure 2: This figure shows the bar plot of the number of assets (out of 92) for which we failed to reject the null hypothesis in the Conditional Coverage (CC) test at various value-at-risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.
Figure 3: This figure shows the bar plot of the number of assets (out of 92) for which we failed to reject the null hypothesis in the Dynamic Quantile (DQ) test at various value-at-risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.

### v6

@@ 8 CRediT author statement
Author 1: Data curation; Formal analysis; Investigation; Validation; Visualization; Writing- Original draft preparation. Author 2: Formal analysis; Investigation; Methodology; Validation; Visualization; Writing- Original draft preparation. Author 3: Conceptualization; Methodology; Project administration; Supervision; Validation; Visualization; Writing- Original draft preparation; Writing- Reviewing and Editing.
Figure 1: The bar plots of number of assets (out of 92) for which we failed to reject the null hypothesis in the Unconditional Coverage (UC) test, at various value-at-risk (VaR) levels, on out-of-sample forecasts for the period January 2015 to September 2023. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.
Figure 2: The bar plot of the number of assets (out of 92) for which we failed to reject the null hypothesis in the Conditional Coverage (CC) test, at various value-at-risk (VaR) levels, on out-of-sample forecasts for the period January 2015 to September 2023. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.
Figure 3: The bar plot of the number of assets (out of 92) for which we failed to reject the null hypothesis in the Dynamic Quantile (DQ) test, at various value-at-risk (VaR) levels, on out-of-sample forecasts for the period January 2015 to September 2023. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.